OpenStudy (anonymous):

An acorn falls from a tree. How fast is the acorn moving after 1.2 seconds?

OpenStudy (anonymous):

So 11.8?

OpenStudy (anonymous):

Yes, your answer is correct. The governing equation is:\[v _{_{f}}=v _{O}+at\]where vf is the final velocity; vO is the starting velocity; a is acceleration; and t is time.

OpenStudy (anonymous):

using eqn. v = u + at thus, v = 0 + 9.8 * 1.2 = 11.76 m/s